Jo Beth left her college town on a bus that traveled 60 mi/h. three hours later, her father left home driving at a speed of 50mi/h to meet the bus. If the college was 345 mi from Jo Beth’s home and the father met the bus as it was arrived at the station, how long did the father drive to meet the bus?
Answer by mananth(16949)

..
let bus station be x miles from home.
..
jo beth traveled 3 hours at 60 mph before father left home
she traveled 180 miles.
..
so she traveled 345-180 = 165 miles from the time her dad started.
time taken by her to travel 165-x miles = 165-x /60 hours
time taken by Dad to reach bus station = x/50
..
times are equal.
x/50 = 165-x / 60
60x=50(165-x)
60x=8250-50x
110x = 8250
x= 8250/110
x= 75 miles distance from home to bus station.